# Cybersecurity threat detection based on a UEBA framework using Deep Autoencoders

### v1

@@ 2.5 UEBA-based anomaly detection framework
The dataset used in this study is derived from multiple real data sources, including Windows events of user activity, emails, and antivirus logs of a financial institution. For this reason, preprocessing plays a crucial role in preparing the raw data for effective anomaly detection. The preprocessing steps involve cleaning, transforming, and encoding the data using the Doc2Vec model described in Section 2.3 to derive features into a format suitable for the anomaly detection model.
The data is ingested as time-series from logs, and is aggregated into fixed windows, summarizing them into key statistics, like total counts and average time intervals. In addition, we perform feature engineering to derive new variables that better capture behavioural patterns, including metrics such as the average time between logins, the ratio of failed to successful logins, and the frequency of antivirus alerts. The set of derived features is detailed in Table 1, with 2 indexing variables and 19 features.
For handling missing values, a lack of activity is assumed to correspond to a zero count, while timing variables (e.g., avg_sec_bet_logins) are imputed using the duration of the aggregation window in seconds as a maximum time.
Meanwhile, text data (specifically, the executable names from processes executed within each window) are combined into a single field (process_list) and encoded using a Doc2Vec model (trained with DBOW) to generate a 64-dimensional embedding vector for each window. These embeddings are concatenated to the derived numerical features, obtaining the final input vector with 83 features.
Finally, all numerical features are normalized using a robust scaler followed by a min-max scaler to ensure that each variable contributes equally to the model. At the end of this preprocessing pipeline, the dataset is clean, structured, and ready for use in the UEBA models and anomaly detection framework.
@@
time Date and time when the data was generated. (used for window aggregation and indexing only) date
@@
Figure 1 illustrates the overall architecture of the proposed UEBA-based anomaly detection framework. The process starts with data collection from multiple sources (e.g., Windows events, emails, antivirus, firewall), which are stored in Splunk Enterprise, a common choice in the industry to collect, index, search, analyse, and visualize large volumes of machine-generated data in real time, as time-series events. The feature extraction pipeline aggregates and processes these events into summary variables and the text data is encoded using Doc2Vec (see Table 1). Once the raw data has been transformed into security events and the features have been computed, this aggregated data is grouped into entities based on business roles (e.g., Customer Management and Executive Positions).
We generated two 1 year historical data sets, one for the Customer Management group (25313 records) and the other for the Executive Positions group (9804 records). Both datasets contain records of user behaviour that have been cleared as normal behaviour by the existing security filters from the institution; however, they may contain a small number of anomalies that evaded these measures. For this reason, we assume the data to be contaminated data (Tian et al., 2023), and thus unlabelled. Both datasets are split into training, validation, and testing sets using a standard split: 20% of the records are held out as the testing set, and the remaining 80% is used for training, of which 20% is further reserved as a validation set. We train a separate deep autoencoder model for each of the user groups. The model architecture starts with 83 input features (including the 19 aggregated features and the 64 components of the Doc2Vec transformation) and compresses them to an 8-dimensional latent space via three hidden layers (with 64, 32, and 16 neurons, respectively). The encoder and decoder networks use ELU activations (except in the first and last layers, which use tanh). Training employs the Adam optimiser, early stopping, and $L_{1}$ regularisation. A decision threshold is determined as the 95th percentile of the reconstruction error on the validation set.
@@ 3 Results and Discussion
We evaluate the proposed UEBA framework in two complementary settings. First, we conduct experiments on real-world data collected from a financial institution on Section 3.1, providing a strong validation of the proposed framework under real operational conditions. Next, we extend these findings on Section 3.2 using simulated data sampled from real attack scenarios, allowing us to systematically probe the system’s response to specific threat vectors and anomalous behaviour of different intensities. By combining these two perspectives –actual enterprise data and controlled simulations– our analysis offers a robust demonstration of how UEBA can enhance security monitoring, highlight anomalous user or entity behaviour, and detect sophisticated cyber threats in an explainable manner.
In the sections that follow, we present the results of these evaluations, focusing on detection rate, residual-space analysis using t-SNE projections, performance under synthetic anomalies, and overall explainability.
@@ 3.1 Performance on real data
Evaluating unsupervised models is challenging in the absence of labelled data in real scenarios. To evaluate the anomaly detection capabilities of the proposed framework, we will assess its ability to learn normal behavioural patterns by analysing the positive rate on the test set. The positive rate is set at 5% based on a decision threshold corresponding to the 95th percentile of the reconstruction error on the validation set. Table 2 shows the positive rates for both user groups (Customer Management and Executive Positions). We can observe how positive rates for both groups are close to the 5% value, indicating that the models are well-calibrated and not overly sensitive to noise in the data. This balance helps reduce the risk of false positives while maintaining a high detection sensitivity.
@@ 3.2 Performance on synthetic anomalies
Figure 3 and 4 presents the detection rate as a function of the anomaly intensity factor $\lambda$ for both user groups (Customer Management and Executive Positions). The results demonstrate that the models reliably detect anomalies when $\lambda>0.7$ for all anomaly types. For specific types, such as login and antivirus anomalies, detection occurs at much lower intensity levels ($\lambda>0.2$). In contrast, process anomalies require higher intensity levels for reliable detection, primarily due to the complexity of encoding text-based data.
@@
At last, we will assess the proposed methodology’s ability to provide explainable model results through the per-feature reconstruction error. Figure 6 shows the logarithm of the per-feature reconstruction error for each model for fully anomalous data ($\lambda=1$). We can observe how that are easier to detect and show this effect more clearly, with higher errors on features related to the anomaly. For instance, the email-related anomalies show the highest reconstruction error on the sent_email_* variables. In addition, for the login anomalies, we get the highest error on the antivirus and login variables.
@@
However, for anomalies that are more challenging, such as process anomalies, the per-variable reconstruction error alone may not suffice in identifying the origin of the anomaly. We speculate that the reason for these anomalies being harder to detect and explain is that they heavily depend on the process encoding by Doc2Vec. However, we have to remark that, when removing the text variables, the detection performance worsened, so the text encoding is providing valuable information for the detection of these anomalies, even if itś not enough for clear interpretation.
We can observe how features related to login patterns, email attachment size, and failed login attempts exhibit the highest contributions to the anomaly score, offering clear indications of potential security incidents. Anomalies in email activity, such as unusually large attachments or an abnormal volume of sent emails, are immediately recognizable. Process-related anomalies, on the other hand, show a more distributed error pattern due to the diverse and complex nature of process command sequences.
These results highlight the strength of the proposed framework in detecting a wide range of anomalies with high accuracy while being explainable. The combination of Deep Autoencoders and Doc2Vec allows for an effective integration of numerical and text-based features, providing a comprehensive view of user behaviour. The ability to visualize residuals and analyse feature-level contributions significantly enhances explainability, making the framework more practical and effective for its use in cybersecurity applications.
@@ 4 Conclusions
This study presented a UEBA-based anomaly detection framework that leverages Deep Autoencoders to identify suspicious activities in a real cybersecurity use-case of a financial institution. By integrating both numerical and text-based features through Doc2Vec embeddings, the proposed approach can capture complex behavioural patterns to detect anomalies that may otherwise remain undetected by more traditional methods. Among the main contributions is a novel theoretical result that proves the equivalence of two common definitions of fully-connected neural network, and a set of experimental evaluations to showcase how advanced deep learning techniques can be employed for explainable, behaviour-based anomaly detection.
@@
Experimental evaluations showed that the the proposed anomaly detection framework achieves a high detection rate, even in challenging conditions with contaminated training data. Additional experiments with synthetic anomalies–reflecting anomalous login patterns, email activity anomalies, and antivirus alerts–confirmed the framework’s robustness and adaptability to diverse cybersecurity threats. A key advantage of our method lies in the ability to perform a residual-based analysis, which enhances explainability by pinpointing specific features that deviate from the normal baseline profile. This explainability is essential in practical cybersecurity contexts where timely responses and a deep understanding of anomalies can significantly improve investigation efficiency and reduce false alarms.
While the proposed framework demonstrated a strong performance, several areas offer opportunities for future improvement. Enhancing text-based feature encoding with more advanced language models (e.g., transformer-based architectures) could improve the detection of process-related anomalies or expand the approach to other text-based sources. Furthermore, integrating explainability techniques like Shapley values and attention mechanisms can provide more insights into feature contributions. Another promising direction is the use of ensemble models to combine multiple anomaly detection algorithms, improving overall performance and reliability.
@@
Conceptualization, J.F, I.O-F, N.M.V and M.S; software, J.F; validation, I.O-F, N.M.V and M.S; formal analysis, J.F, I.O-F, N.M.V and M.S; investigation, J.F; resources, I.O-F, M.S; data curation, J.F; writing-original draft, J.F, I.O-F, N.M.V; writing-review & editing, I.O-F, M.S, and N.M.V; visualization, J.F; supervision, I.O-F, M.S, and N.M.V; project administration, I.O-F; funding acquisition, I.O-F, M.S.

### v2

@@ 2.5 UEBA-based anomaly detection framework
The dataset used in this study is derived from multiple real data sources, including Windows events of user activity, emails, and antivirus logs from a financial institution. For this reason, preprocessing plays a crucial role in preparing the raw data for effective anomaly detection. The preprocessing steps involve cleaning, transforming, and encoding the data using the Doc2Vec model described in Section 2.3 to derive features into a format suitable for the anomaly detection model.
The data is ingested as time-series from logs and is aggregated into fixed windows, summarising them into key statistics, like total counts and average time intervals. In addition, we perform feature engineering to derive new variables that better capture behavioural patterns, including metrics such as the average time between logins, the ratio of failed to successful logins, and the frequency of antivirus alerts. The set of derived features is detailed in Table 2, with 2 indexing variables and 19 features.
For handling missing values, a lack of activity is assumed to correspond to a zero count, while timing variables (e.g., avg_sec_bet_logins) are imputed using the duration of the aggregation window in seconds as a maximum time. Meanwhile, text data (specifically, the executable names from processes executed within each window) are combined into a single field (process_list) and encoded using a Doc2Vec model (trained with DBOW) to generate a 64-dimensional embedding vector for each window. These embeddings are concatenated to the derived numerical features, obtaining the final input vector with 83 features.
Finally, all numerical features are normalised using a robust scaler followed by a min-max scaler to ensure that each variable contributes equally to the model. At the end of this preprocessing pipeline, the dataset is clean, structured, and ready to be used in the UEBA models and anomaly detection framework.
@@
time Date and time when the data was generated (used for window aggregation and indexing only) date
@@
Figure 1 illustrates the overall architecture of the proposed UEBA-based anomaly detection framework, where the numbered arrows correspond to the different stages of the process (1–5). The process starts with data collection (1) from multiple sources (e.g., Windows events, emails, antivirus, firewall), which are stored in Splunk Enterprise, a common choice in the industry to collect, index, search, analyse, and visualise large volumes of machine-generated data in real time, as time-series events. The feature extraction pipeline aggregates and processes these events into summary variables, while the text data is encoded using Doc2Vec (see Table 2). Once the raw data has been transformed into security events and the features have been computed, this aggregated data is grouped into entities based on business roles, e.g. Customer (CM) and Executive Positions (EP) and stored at the UEBA index in Splunk (2). This aggregated UEBA dataset can now be used to train the Autoencoder and Doc2Vec models (see Algorithm 1), which are later stored in a model store (3). The hyperparameters used to train both models are available in Table 3, and the network diagram is presented in Figure 2. Once the models are trained and stored, they are made available through the model store (4) to generate predictions for new data (see Algorithm 2) coming from the index (4) in near real-time, including the residual-based anomaly scores which are stored in the alert databased, and therefore available for plotting and advanced analysis (5).
@@
Table 3: Hyperparameters of the UEBA anomaly detection framework for the Customer Management (CM) and Executive Positions(EP) roles
Component Setting
Doc2Vec 64-dimensional embeddings; DBOW architecture; window size = 5; epochs = 20
Autoencoder Input size = 83 features; hidden layers = [64, 32, 16, 8, 16 32, 64]; latent dimension = 8; activations = ELU (internal), tanh (first/last)
Training Optimiser = Adam (lr = 0.001 (CM) / 0.01 (EP) [^1] The higher learning rate in the EP role was selected via grid search in combination with the larger batch size, resulting in stable convergence and improved generalisation.); batch size = 64 (CM) / 256 (EP); early stopping (patience = 10, monitor = validation MSE); $L_{1}$ regularization ($\lambda=0.001$); validation split = 20%
Figure 2: Network diagram of the proposed autoencoder.
Input: Windowed log events $\{\mathcal{E}_{i}\}$
Output: Trained autoencoder $f_{\theta}$, Doc2Vec encoder $\phi$, scaler, decision threshold $\tau$
  1. 1.
    Extract numerical features $n_{i}=g_{\text{num}}(\mathcal{E}_{i})$.
  2. 2.
    Encode process list tokens $\mathcal{T}_{i}$ with Doc2Vec (DBOW) to obtain $t_{i}=\phi(\mathcal{T}_{i})$.
  3. 3.
    Concatenate: $x_{i}=[n_{i};t_{i}]$, then apply scaling to get $\tilde{x}_{i}$.
  4. 4.
    Train deep autoencoder $f_{\theta}$ to minimize reconstruction loss with $L_{1}$ regularization and early stopping.
  5. 5.
    Compute scores $s_{i}=\|\tilde{x}_{i}-f_{\theta}(\tilde{x}_{i})\|_{1}$ on validation data.
  6. 6.
    Set threshold $\tau$ as the 95th percentile of $\{s_{i}\}$.
Algorithm 1 Training of the UEBA framework
Input: New window $\mathcal{E}$; trained $f_{\theta}$, $\phi$, scaler; threshold $\tau$
Output: Score $s$, residual vector $r$, decision $y\in\{\text{normal},\text{anomaly}\}$
  1. 1.
    Extract features: $n=g_{\text{num}}(\mathcal{E})$, $t=\phi(\mathcal{T})$, $x=[n;t]$.
  2. 2.
    Scale input: $\tilde{x}=\mathrm{scale}(x)$.
  3. 3.
[…]
  4. 4.
    Compute residuals $r=\tilde{x}-\hat{x}$ and anomaly score $s=\|r\|_{1}$.
  5. 5.
    Decision: $y=\mathbb{I}[s\geq\tau]$. Return $(s,r,y)$ for alerting and visualisation.
Algorithm 2 Inference and anomaly scoring
We generated two different datasets of one year of historical data, one for the Customer Management group (25313 records) and the other for the Executive Positions group (9804 records). Both datasets contain records of user behaviour that have been cleared as normal behaviour by the existing security filters from the institution; however, they may contain a small number of anomalies that evaded these measures. For this reason, we assume the data to be contaminated data [34], and thus unlabelled. Both datasets are split into training, validation, and testing sets using a standard split: 20% of the records are held out as the testing set, and the remaining 80% is used for training, of which 20% is further reserved as a validation set. We train a separate deep autoencoder model for each of the user groups. The model architecture starts with 83 input features (including the 19 aggregated features and the 64 components of the Doc2Vec transformation) and compresses them to an 8-dimensional latent space via three hidden layers (with 64, 32, and 16 neurons, respectively). The encoder and decoder networks use ELU activations (except in the first and last layers, which use tanh). Training employs the Adam optimiser, early stopping, and $L_{1}$ regularisation. A decision threshold $\tau$ is determined as the 95th percentile of the reconstruction error on the validation set.
@@ 3 Results and Discussion
We evaluate the proposed UEBA framework in two complementary settings. First, we conduct experiments on real-world data collected from a financial institution in Section 3.1, providing a strong validation of the proposed framework under real operational conditions. Next, we extend these findings in Section 3.2 using simulated data sampled from real attack scenarios, allowing us to systematically probe the system’s response to specific threat vectors and anomalous behaviour of different intensities. By combining these two perspectives–actual enterprise data and controlled simulations—our analysis offers a robust demonstration of how UEBA can enhance security monitoring, highlight anomalous user or entity behaviour, and detect sophisticated cyber threats in an explainable manner.
In the following sections, we present the results of these evaluations, focusing on detection rate, residual-space analysis using t-SNE projections, performance under synthetic anomalies, and overall explainability.
@@ 3.1 Performance on real data
Evaluating unsupervised models is challenging in the absence of labelled data in real scenarios. To evaluate the anomaly detection capabilities of the proposed framework, we assess its ability to learn normal behavioural patterns by analysing the positive rate on the test set. The decision threshold $\tau$ is fixed at the 95th percentile of the reconstruction error on the validation set. Table 4 shows the positive rates for both user groups (Customer Management and Executive Positions). We can observe how positive rates for both groups are close to the 5% value, indicating that the models are well-calibrated. While the overall calibration is consistent across groups, we note that the positive rate of the Executive Positions model (4.61%) is slightly below the 5% target, whereas the Customer Management model (5.09%) is slightly above it. This difference is minor in absolute terms, but it reflects the fact that user groups with fewer samples and more heterogeneous activity patterns may lead to tighter thresholds and a slightly more conservative model. Such variations are expected in UEBA applications, where behaviour differs by role, and they showcase the importance of tailoring models to specific roles or groups.
These positive-rate results constitute the primary evaluation of our framework in an unsupervised setting. Since the threshold was fixed during training, the fact that calibration holds on the test set shows that the models generalise well without relying on labelled anomalies. Standard supervised metrics such as precision, recall, or F1 cannot be meaningfully computed here, since the dataset lacks anomaly labels, and the data is taken as possibly contaminated data. In addition, as reported in Section 3.2, we were provided with a small set of attack events. Although these real anomalies were valuable, their very limited number made the evaluation less informative: the models detected them almost trivially, with a true positive rate close to $100\%$. This limitation motivated the use of synthetic anomaly experiments, which should be regarded as a complementary stress test probing robustness under varying anomaly intensity rather than as the main evaluation.
@@ 3.2 Performance on synthetic anomalies
This convex interpolation scheme provides a simple yet controlled way to validate the anomaly detection sensitivity. By varying the anomaly intensity factor $\lambda$, we can gradually shift normal samples toward real anomalies and observe how the model responds, while ensuring that the resulting points remain inside the convex hull of the observed feature space, reducing the risk of generating unrealistic samples. At the same time, this construction mainly alters feature magnitudes and does not capture temporal dependencies, cross-feature correlations, or coordinated attack strategies. Therefore, synthetic anomalies should be regarded as a complementary stress test of robustness, rather than a substitute for larger-scale evaluation on real-world attack data.
@@
Figure 4 and 5 present the detection rate as a function of the anomaly intensity factor $\lambda$ for both user groups (Customer Management and Executive Positions). The results demonstrate that the models reliably detect anomalies when $\lambda>0.7$ for all anomaly types. For specific types, such as login and antivirus anomalies, detection occurs at much lower intensity levels ($\lambda>0.2$). In contrast, process anomalies require higher intensity levels for reliable detection, primarily due to the complexity of encoding text-based data. While Doc2Vec provides a compact and efficient encoding, its averaging nature can obscure rare but highly informative tokens. More expressive embeddings could, in principle, capture richer temporal and syntactic structure, though at higher computational cost (see Section 4 for possible alternatives for future work).
These detection-vs-intensity curves serve as surrogate evaluation metrics in the absence of extensive labelled examples. Rather than relying on precision or recall, which cannot be meaningfully computed here, we evaluate how detection rates evolve as normal samples are gradually shifted toward known attack behaviours. This provides a principled and interpretable way to assess sensitivity under scarce-label conditions, showing that the framework can reliably separate normal from anomalous behaviour once the anomalies reach sufficient intensity. In this sense, the detection–vs–intensity curves can be interpreted as an analogue of statistical power curves: the anomaly intensity parameter $\lambda$ plays a role similar to the effect of the sample size, and the resulting curves quantify how quickly the detector achieves high detection probability as anomalies become more pronounced.
@@
At last, we will assess the proposed methodology’s ability to provide explainable model results through the per-feature reconstruction error. Figure 7 shows the logarithm of the per-feature reconstruction error for each model for fully anomalous data ($\lambda=1$). We can observe that they are easier to detect, and this effect becomes more evident, with higher errors appearing on features related to the anomaly. For instance, the email-related anomalies show the highest reconstruction error on the sent_email_* variables. In addition, in the case of login anomalies, the highest errors are observed in the antivirus and login variables.
However, for anomalies that are more challenging, such as process anomalies, the per-variable reconstruction error alone may not suffice in identifying the origin of the anomaly. We speculate that the reason these anomalies are harder to detect and explain is that they heavily depend on the process encoding by Doc2Vec. However, we have to remark that, when removing the text variables, the detection performance worsened, so the text encoding provides valuable information for the detection of these anomalies, even if itś not enough for clear interpretation.
@@
We can observe how features related to login patterns, email attachment size, and failed login attempts exhibit the highest contributions to the anomaly score, offering clear indications of potential security incidents. Anomalies in email activity, such as unusually large attachments or an abnormal volume of sent emails, are immediately recognisable. Process-related anomalies, on the other hand, show a more distributed error pattern due to the diverse and complex nature of process command sequences.
These results highlight the strength of the proposed framework in detecting a wide range of anomalies with high accuracy while being explainable. The combination of Deep Autoencoders and Doc2Vec allows for an effective integration of numerical and text-based features, providing a comprehensive view of user behaviour. The ability to visualise residuals and analyse feature-level contributions significantly enhances explainability, making the framework more practical and effective for its use in cybersecurity applications.
@@ 4 Conclusions
This study presents a UEBA-based anomaly detection framework that leverages Deep Autoencoders to identify suspicious activities within a real-world cybersecurity use case in a financial institution. By integrating both numerical and text-based features through Doc2Vec embeddings, the proposed approach can capture complex behavioural patterns to detect anomalies that may otherwise remain undetected by more traditional methods. A key contribution is a novel theoretical result proving the equivalence of two common definitions of fully connected neural networks, thereby grounding our autoencoder design within universal approximation theory. In addition, we provide a set of experimental evaluations to showcase how advanced deep learning techniques can be employed for explainable, behaviour-based anomaly detection.
@@
Experimental evaluations showed that the proposed anomaly detection framework achieves a high detection rate, even in challenging conditions with contaminated training data. Additional experiments with synthetic anomalies–reflecting anomalous login patterns, email activity anomalies, and antivirus alerts–confirmed the framework’s robustness and adaptability to diverse cybersecurity threats. A key advantage of our method lies in the ability to perform a residual-based analysis, which enhances explainability by pinpointing specific features that deviate from the normal baseline profile. This explainability is essential in practical cybersecurity contexts where timely responses and a deep understanding of anomalies can significantly improve investigation efficiency and reduce false alarms.
Compared to more traditional anomaly detection techniques such as PCA-based monitoring, statistical thresholds, or clustering approaches, our framework offers several distinctive advantages. As also summarised in Table 1, classical methods typically assume clean or labelled training data, can only be applied to numerical features, and provide limited interpretability. In contrast, our approach is explicitly designed to operate under data contamination, integrates both numerical and textual features through Doc2Vec, and leverages residual-based scores to deliver ante-hoc explanations. These aspects underline the novelty and uniqueness of our framework in enterprise UEBA scenarios.
While the proposed framework demonstrated strong performance, several limitations should be acknowledged. First, the amount of labelled anomalies was extremely limited, which prevented the use of standard supervised metrics and restricted baseline comparisons. Second, the reliance on Doc2Vec embeddings for process features, while efficient, may obscure rare but highly informative tokens. Finally, the evaluation was performed on sensitive institution-specific data that cannot be released, which constrains reproducibility.
These limitations point to several directions for future research. A deeper and systematic comparison with standard unsupervised baselines (e.g., Isolation Forest, LOF, one-class SVM) remains a priority, ideally using a public benchmark dataset such as the CERT Insider Threat dataset [35] or datasets with more textual features. Future work should also explore more expressive text encoders, such as transformer-based or recurrent architectures, especially for longer and more natural textual features (e.g., email content, web addresses). In parallel, advancing explainability remains essential. Our residual-based approach already highlights which features are most difficult to reconstruct, offering ante-hoc insights into the features that most influence anomaly detection. Post-hoc techniques such as Shapley values [36] or LIME [37] could complement this by accounting more explicitly for feature interactions and correlations, providing analysts with a second layer of explanation to validate and enrich our residual-based insights. Moreover, counterfactual analyses [38] further extend this perspective by showing what minimal behavioural changes would render an anomalous user or entity normal, which could help analysts perform root-cause analysis and guide remediation strategies. Finally, ensemble strategies that combine complementary anomaly detectors offer a promising route to enhance robustness and reliability in practical deployments.
Data Availability
The real world enterprise logs analysed in this study were provided by a financial institution under confidentiality and security agreements and are not publicly available. Aggregate statistics and the synthetic anomaly generation procedure are described in Section 2 and Subsection 3.2.
@@
Conceptualization, J.F, I.O-F, N.M.V and M.S; software, J.F; validation, I.O-F, N.M.V and M.S; formal analysis, J.F, I.O-F, N.M.V and M.S; investigation, J.F; resources, I.O-F, M.S; data curation, J.F; writing—original draft, J.F, I.O-F, N.M.V; writing—review & editing, I.O-F, M.S, and N.M.V; visualization, J.F; supervision, I.O-F, M.S, and N.M.V; project administration, I.O-F; funding acquisition, I.O-F, M.S.